OpenStudy (anonymous):

Find the value of 7 times x squared all over the square root of x to the fifth power. 7 times the square root of x, all over x. 7 over x. 7 times the square root of x. 7 times x all over the square root of x.

OpenStudy (anonymous):

Does it look like this \[\frac{ 7 \times x ^{2} }{ \sqrt{x ^{5}} }\]

OpenStudy (anonymous):

no times between 7 and x2

OpenStudy (anonymous):

You can rewrite that expression as \[\frac{ 7 }{ x ^{\frac{ 5 }{ 2 }-2} }\]

OpenStudy (anonymous):

Do you know what \[\frac{ 5 }{ 2 } - 2 =\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about this  can you find what this equals to ?

OpenStudy (anonymous):

half 1/2

OpenStudy (anonymous):

good... so now we know that the expression looks like \[\frac{ 7 }{ \sqrt{x} }\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

We can simplify this radical to make it look likes this..\[\frac{ 7 }{ \sqrt{x} } \times \frac{ \sqrt{x} }{ \sqrt{x} } = \frac{ 7\sqrt{x} }{ x }\]

OpenStudy (anonymous):

can you find a sentence from your choices that reflect this mathematical statement..

OpenStudy (anonymous):

that the answer

OpenStudy (anonymous):

really ! well done =))